OpenStudy (anonymous):

Clear fractions or decimals, solve, and check. 5/-6 + x= 1/-2 2/-3

OpenStudy (callisto):

\[\frac{5}{-6}+x = \frac{1}{-2} \frac{2}{-3}\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Hmm... for the right side, not addition / subtraction sign between the 2 fractions?

OpenStudy (anonymous):

what you thinking. I been looking at this for the longs time

OpenStudy (callisto):

Hmm.. if there're no signs between the 2 fractions, then it would be multiplication. \[\frac{5}{-6}+x = \frac{1}{-2}\frac{2}{-3}\]\[\frac{5}{-6}+x = \frac{1(2)}{-2(-3)}\]\[\frac{5}{-6}+x = \frac{2}{6}\]\[x = \frac{2}{6}+ \frac{5}{6} = \frac{7}{6}\]

OpenStudy (callisto):

But I don't know why, I think there is a sign between the 2 fractions, like this \[\frac{5}{-6}+x=\frac{1}{-2}+\frac{2}{-3}\]

OpenStudy (anonymous):

l look again in the book just to be safe

OpenStudy (callisto):

SURE!!!~

OpenStudy (anonymous):

no sign between the 2 fractions

OpenStudy (callisto):

Then, that's it :)

OpenStudy (anonymous):

thank you :)

OpenStudy (callisto):

Welcome :)

OpenStudy (anonymous):

I'm write it down want to help with just 3 more

OpenStudy (anonymous):

you same good at it and you make it easy for me

OpenStudy (callisto):

I would help if I could :)